The equation of the normal at t=pie/2 to the curve x=2sint,y=2cost is

Grade:11

1 Answers

Aditya Gupta
4 years ago
hello kavitha, note that at t=pie/2, x=2sint= 2 and y=2cost= 0
now, dy/dx at t=pie/2= (dy/dt)/(dx/dt)= – 2sint/2cost= – tant = not defined at t=pie/2.
which means that there is a vertical tangent at t=pie/2. since normal is perpendicular to tangent, hence it will hve to be horizontal, with slope= 0
so, eqn of normal: y= 0 or the x axis 
kindly approve :)